Given that k =[2 1]




[3 4]




Find the matrix k2+k+i,where i is the (2× 2) identify matrix






1
Expert's answer
2022-01-31T12:55:42-0500
"K^2=\begin{bmatrix}
 2 & 1 \\
 3 & 4
\end{bmatrix}\begin{bmatrix}
 2 & 1 \\
 3 & 4
\end{bmatrix}"

"=\begin{bmatrix}
 2(2)+1(3) & 2(1)+1(4) \\
 3(2)+4(3) & 3(1)+4(4)
\end{bmatrix}=\begin{bmatrix}
 7 & 6 \\
 18 & 19
\end{bmatrix}"

"K^2+K+I=\begin{bmatrix}
 7 & 6 \\
 18 & 19
\end{bmatrix}+\begin{bmatrix}
 2 & 1 \\
 3 & 4
\end{bmatrix}+\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

"=\begin{bmatrix}
 7+2+1 & 6+1+0 \\
 18+3+0 & 19+4+1
\end{bmatrix}=\begin{bmatrix}
 10 & 7 \\
 21 & 24
\end{bmatrix}"

"K^2+2K+I=\begin{bmatrix}
 10 & 7 \\
 21 & 24
\end{bmatrix}"